A tailor must make 8 of the same size suits for a client. The first suit took 13.3 hours to make and the last suit took 8.8 hours to make. What was the tailor's learning curve % of this run of suits assuming a steady state of production is never reached? Express your answer in % to the nearest 0.1%
Answer #1

Tn=T1*n^b

T1=13.3 hours

T8=8.8 hours

T8=T1*8^b

8.8=13.3*8^b

2/3=8^b

b=lnp/ln2 where p is learning rate

0.67=8^(lnp/ln2)

ln(0.67)=lnp*ln8/ln2=ln(p^3)

0.67=p^3

p=87.5%

Therefore we have learning rate of 87.5 %